# Generating Attribute-Aware Human Motions from Textual Prompt

### v1

@@ Abstract
Text-driven human motion generation has recently attracted considerable attention, allowing models to generate human motions based on textual descriptions. However, current methods neglect the influence of human attributes—such as age, gender, weight, and height—which are key factors shaping human motion patterns. This work represents a pilot exploration for bridging this gap. We conceptualize each motion as comprising both attribute information and action semantics, where textual descriptions align exclusively with action semantics. To achieve this, a new framework inspired by Structural Causal Models is proposed to decouple action semantics from human attributes, enabling text-to-semantics prediction and attribute-controlled generation. The resulting model is capable of generating realistic, attribute-aware motion aligned with the user’s text and attribute inputs. For evaluation, we introduce HumanAttr, a comprehensive dataset containing attribute annotations for text-motion pairs, setting the first benchmark for attribute-aware text-to-motion generation. Extensive experiments on the new dataset validate our model’s effectiveness.
1 Introduction
Figure 1: Data samples from HumanAttr dataset with textual descriptions and annotated human attributes. Notice that the motion patterns of subjects with different attributes vary significantly.
@@ 1 Introduction
In recent years, text-driven human motion generation has gained significant attention [92, 79, 80, 120, 128, 24, 123, 124, 13, 93, 117, 99, 41] in both research community and industry owing to its potential applications in video games, filmmaking, virtual reality, and robotics. The goal is to generate realistic human motion sequences based on given textual descriptions. However, human motion patterns are significantly influenced by human attributes such as age, gender, weight, and height. As illustrated in Figure 1, individuals with different attributes demonstrate distinct motion patterns. For instance, the movements of an elderly person and a teenager differ substantially. Therefore, incorporating human attributes into motion generation is crucial for producing realistic and context-aware movements. Also, users may need to generate motions that correspond to specific human attributes, demanding text-to-motion generation controlled by attributes. Despite this, none of the existing methods have addressed this issue.
@@
Incorporating human attributes into text-driven motion generation faces two major challenges. The first is that a motion sequence consists of both action semantics (e.g., walking, running) and human attributes (e.g., gender, age), whereas textual descriptions usually focus solely on the action semantics aspect. Existing text-to-motion frameworks [92, 79, 80, 120, 128, 71] typically align text and motion in a shared space without distinguishing between semantic content and human attribute information, which may hinder the alignment. For example, MotionCLIP [92] aligns textual and motion features in the CLIP latent space, while MotionGPT [42] obtains a unified token vocabulary for text and motion. This necessitates a new framework that can separate action semantics from human attributes, ensuring that texts are accurately aligned with semantic content.
@@
The second challenge lies in the absence of a large-scale text-to-motion dataset that includes annotations for a broad spectrum of human attributes. Existing datasets either lack human attribute annotations entirely [24, 66, 82, 65, 11], encompass subjects with only a small range of distinct attributes [81, 40, 95, 73, 50], or have very limited scale [1, 55]. For example, the KIT dataset [81] includes attribute annotations for 4k motion sequences, but 90% of the 55 subjects are aged between 18 and 45. Similarly, the ETRI-LivingLab dataset [50] features 50 subjects, all aged between 64 and 80. This data gap impedes the development of attribute-aware motion generation models capable of generalizing to diverse textual descriptions and human attributes.
@@
This work represents a pilot exploration that incorporates human attributes to generate high-quality 3D human motion aligned with prompted text and attribute controls. To address the above-mentioned semantics-attribute coupling in human motions, we introduce AttrMoGen (Attribute-controlled Motion Generator). It comprises a Semantic-Attribute Decoupling VQVAE (Decoup-VQVAE) that extracts attribute-free semantic tokens from human motions, and a Semantics Generative Transformer that predicts these semantic tokens from textual prompts. To eliminate attribute information from the semantic tokens, we leverage Structural Causal Model (SCM) [77], formulating the problem as causal factor disentanglement. In particular, action semantics $S$ are causal factors determining target tokens $Y$, while human attributes $A$ are non-causal for $Y$ but crucial for constructing raw motion $X$. The causal model is trained using a causal information bottleneck, which disentangles $S$ from $A$ by limiting their mutual information $I(S;A)$. Additionally, a bottleneck term $-\lambda I(X;S)$ is used to restrict information flow from $X$ to $S$, ensuring that $S$ captures only the essential semantics. During inference, the Semantics Generative Transformer predicts semantic tokens from text input, which are combined with attributes input to generate attribute-aware motion via the Decoup-VQVAE decoder.
@@
We compiled a comprehensive dataset called HumanAttr, which includes texts and attribute annotations for each motion sequence. HumanAttr integrates data from multiple sources, resulting in a collection of 18.2k motion sequences from 640 subjects with a wide range of attributes. Example data entries are presented in Figure 1. Each attribute label consists of the subject’s age and gender. A portion of the data ($\sim$ 74%) also contains weight and height information, but lacks age diversity thus are not utilized in the main experiments. An additional experiment on this subset that contains weight and height is included in supplementary materials for reference. Extensive experiments are conducted on the dataset, evaluating a range of current representative text-to-motion models, including our proposed AttrMoGen model. The results demonstrate its ability to generate realistic human motion that aligns well with specified texts and human attributes.
2 Related Work
Human Motion Generation. Generating realistic human motion that aligns with specific control signals has been a longstanding problem. A prominent focus is text-to-motion generation. Early works [3, 21, 92, 79, 80, 115, 83] primarily focused on establishing a joint embedding space for motion sequences and texts. For example, MotionCLIP [92] aligns human motion embeddings with the CLIP latent space using cosine similarity loss.
Recent development has been dominated by autoregressive models and diffusion models. The former [52, 131, 120, 128, 24, 116, 136, 38, 71, 135] represent motions with discrete tokens and implement text-to-motion generation via next token prediction. For instance, T2M-GPT [121] employs a VQ-VAE to encode motion into discrete tokens, which can be generated autoregressively conditioned on text tokens. Among diffusion models and their variants [44, 118, 127, 6, 101, 47, 12, 111, 30, 31, 34, 108, 107, 133, 19, 45, 85, 64, 37, 90, 14, 130], MotionDiffuse [123] introduces a DDPM-based architecture that enables realistic motion synthesis with part-level manipulability. Besides texts, motion can also be generated conditioned on other multimodal inputs, including action labels [23, 78, 8, 109, 54, 18], audio [57, 56, 88, 96, 59, 60, 134, 22, 119, 34, 67, 114], history or future poses [4, 51, 32, 48], human trajectories [86, 46, 107, 104, 98, 35, 62], 3D-scenes or objects [7, 105, 9, 110, 53, 17, 112, 10], etc.
Recently, some studies [106, 132, 28, 29, 122, 49, 89] explore style-based motion models, emphasizing how an action is performed (e.g., proud, depressed, angry). In contrast, attributes denote people’s inherent and permanent characteristics (e.g., gender and age) which directly affect the biomechanics of their motion. Although some style labels like ”childlike” may appear attribute-related, they typically depict performers are imitating specific motion patterns. The distinction is evident in widely-used style-motion datasets [2, 76] where one single actor, characterized by certain attribute profile, performs actions in various styles.
Text-motion Datasets. With the proliferation of mocap-based databases [39, 15, 87, 69, 75, 74], numerous text-motion datasets have been proposed. Conventional ones like KIT [81] and HumanML3D [24] provide motion with corresponding sequence-level textual descriptions, primarily developed for text-to-motion task. Subsequent works have advanced the field by incorporating multi-person scenarios [65, 20, 102], addressing long motion sequences [61, 30], enhancing the quality of text descriptions [91, 11], adding full-body text descriptions [66], and providing frame-level or segment-level annotations [82, 100].
In addition to datasets with solely text annotations, recent studies introduced diverse annotations’ modalities, such as object or 3D scene [103, 129, 43], music [126, 58] and speech [126, 113, 68]. Several motion datasets [1, 81, 95, 73, 40, 50] provide the attribute information of each subject, yet are limited in data scale or attribute range. For instance, 90% of the subjects in KIT [81] are aged between 18 and 45. This indicates the lack of large-scale collections that encompass a wide range of human attributes.
@@ 2 Related Work
BMLmovi [95] 86 1,801 161.8 [17, 33]
EKUT [81] 4 348 30.1 [24, 44]
ETRI-Activity3D [40] 100 3,727 691.8 [21, 88]
ETRI-LivingLab [50] 50 600 160.5 [64, 80]
K3Da [55] 54 236 34.5 [18, 81]
Kinder-Gator [5] 20 362 34.9 [5, 32]
KIT [81] 55 4,231 463.1 [15, 55]
Nymeria [73] 264 6,850 552.4 [18, 50]
SFU [1] 7 44 6.4 [18, 30]
@@
Table 1: Statistics of the collected sub-datasets. Note that for  [95, 81, 1], only entries with HumanML3D [25] text annotations are counted. The statistics are gathered after data filtering.
3 The HumanAttr Dataset
Dataset Collection. Existing benchmark datasets for human motion generation such as HumanML3D [24] and KIT [81] primarily consist of motion sequences and corresponding text descriptions, lacking the annotations for human attributes. To address this gap, we compile a comprehensive dataset that includes data entries of (motion, text, attribute). The attribute labels contain age and gender for each data entry, and a portion of the data ($\sim$74%) also have weight and height information. The dataset, named HumanAttr, integrates data from various sources, annotated with diverse human attributes and textual descriptions. As illustrated in Table 1, HumanAttr consists of nine sub-datasets that encompass a variety of subjects with varying attributes. In total, it contains 18.2k motion sequences from 640 subjects, with ages ranging from 5 to 88 years and balanced gender. Figure 7 presents the distribution of age, gender, motion duration, and text length within the dataset.
Dataset Processing. This section provides a brief introduction to the pipeline for processing HumanAttr. Initially, data from different sources are converted into a 3D coordinates representation of $L\times V\times 3$, where $L$ denotes the sequence length and $V$ is the number of markers. Since the number of markers varies across different data sources, we adopt the method from [63] and use gradient descent optimization to fit the marker data into unified parameterized SMPL model [70]. To make the dataset compatible with previous work in the field, processing protocol from HumanML3D [24] is then applied, resulting in a 263-dimensional input vector.
@@ 3 The HumanAttr Dataset
However, some of the motion data exhibited serious jittering issues caused by the Kinect system. To address this, additional processing is applied to specific sub-datasets [40, 50, 55, 5]. Motions with excessive jitter or outlier coordinates are discarded, and a 1-D Gaussian denoising is applied to enhance motion quality. The dataset is further augmented by mirroring the motion data following the approach in HumanML3D to increase the training samples and improve model generalization. For text annotations, all sub-datasets include either textual descriptions or action category labels. Action labels are expanded into full text descriptions using the information from the original database, ensuring each motion entry has a meaningful text.
4 Our Proposed Method
To tackle this challenge, we propose an architecture composed of two main components. The overall architecture is illustrated in Figure 3. The first component is a Semantic-attribute Decoupling VQVAE (Decoup-VQVAE), inspired by Structural Causal Model (SCM) [77]. Specifically, the encoder removes attribute information from the raw motion with the help of a causal information bottleneck, resulting in attribute-free semantic tokens via vector quantization. The decoder can then reconstruct the original motion using both the semantic tokens and attribute labels. The second component is a Semantics Generative Transformer, which predicts semantic tokens from text by a generative transformer. During inference, semantic tokens are generated from textual input and combined with user-defined attributes to generate attribute-aware human motion.
4.1 Semantic-attribute Decoupling VQVAE
To obtain attribute-free semantic tokens, we employed Structural Causal Model (SCM) to formulate the problem as causal factor disentanglement. Specifically, let $X$ represent the raw motion and $Y$ represent the desired semantic tokens. The action semantics $S$ are the causal factors that determine $Y$. Human attributes, denoted as $A$, are non-causal factors that are independent of $Y$, but essential in constructing the raw motion $X$. This leads to our causal formulation, with the SCM illustrated in Figure 4. Our aim is to learn an encoder $S=f(X)$ that can decouple the $Y$-causative action semantics $S$ from raw motion $X$, and a decoder $\hat{X}=g(S,A)$ that can reconstruct the motion $X$ from $S,A$. Ideally, $f$ should clearly separate $S$ from $A$, retaining only the essential semantic information of the raw motion. Inspired by [125, 36], we introduce a causal information bottleneck (CIB) to tackle this challenge. Formally, The CIB objective function is defined as
@@ 4.1 Semantic-attribute Decoupling VQVAE
where $I(;)$ denotes the mutual information and $\lambda$ is the bottleneck weight. [125] proves that maximizing the CIB objective is equivalent to maximizing the mutual information $I(X,Y;S,A)$ between the latent factors $S$ and $A$ and the variables $X$ and $Y$ with an information bottleneck. The role of each term and our customized implementation will be demonstrated below.
Figure 4: Structural Causal Model (SCM) for our Decoup-VQVAE. Our objective is to learn an encoder capable of decoupling the $Y$-causative action semantics $S$ from raw motion $X$.
@@
The term $-I(S;A)$ restricts the mutual information between $S$ and $A$, which plays a crucial role in disentangling the causal factor $S$ from the non-causal factor $A$. In our case, note that the mutual information has the following upper bound:
@@
The information bottleneck term $-\lambda I(X;S)$ limits the information flow from $X$ to $S$, ensuring that $S$ contains only essential semantics without excessive irrelevant information, fostering a robust and compact representation [94]. The mutual information $I(X;S)$ can be measured using the Kullback-Leibler (KL) divergence:
@@
where $A^{-}$ is the counterfactual attribute obtained by randomized the original attribute $A$. Suppose $\boldsymbol{X}$ is a batch of samples and $\boldsymbol{X}^{-}$ is their corresponding counterfactuals. As the encoder $S=f(X)$ serves as an estimator for the distribution $p(S|X)$, the KL-divergence in Equation 10 can be minimized by aligning $f(\boldsymbol{X})$ and $f(\boldsymbol{X}^{-})$. To achieve this, we define the following similarity matrix analogous to [72]:
@@
The remaining terms, $I(X;S,A)+I(Y;S)$, ensure that the information contained in $S$ and $A$ are sufficient to reconstruct $X$, while also guarantee that $S$ is informative enough to deduce $Y$. Note the following lower bound derived by [125]:
@@
The second term $I(Y;S)$ ensures the deduction from the causal factor $S$ to the target semantic token $Y$, which is, by definition, the quantization process from the semantic embedding to the codebook index. Note that $I(Y;S)=H(Y)-H(Y|S)$, where $H(Y)$ reflects the full exploitation of the codebook and $-H(Y|S)$ demands precise indexing from $S$ to $Y$. This can be addressed by the loss $\mathcal{L}_{commit}+\mathcal{L}_{embed}$ from VQVAE [97]. In summary, the overall loss function is formulated as
@@
4.2 Semantics Generative Transformer
To predict semantic tokens from text, a Semantics Generative Transformer is employed, whose architecture can be either a GPT-like autoregressive model [121, 128, 71] or a masked transformer [27]. We adopt MoMask [27] due to its strong performance and efficiency. During training, the model aims to predict randomly masked semantic tokens conditioned on the CLIP [84] features of the textual input, following a BERT-like scheme [16]. At inference time, semantic tokens are first predicted from the text using the Semantics Generative Transformer. These tokens, combined with attribute inputs (e.g., age, gender), are then used to generate the motion sequence via the Decoup-VQVAE decoder. Figure 3 presents a schematic illustration.
5 Experiments
Implementation and Evaluation Details.
The dataset is split into training, validation and testing sets in an 80%, 5%, and 15% ratio. All data preprocessing protocols follow those used in HumanML3D [25], and the input dimension for the encoder is 263. For the model architecture, MoMask [27] is adopted due to its superior performance and efficiency. This architecture consists of a residual VQVAE, a masked transformer, and a residual transformer. The network architecture for the proxy classifier $h$ is the same as VQVAE encoder.
For attribute control, we categorize the attributes into discrete labels to accommodate sub-datasets that only provide age groups (e.g., Nymeria) and enable in-depth analyses based on categorization. To be specific, gender is labeled as 0 for male and 1 for female, while age is divided into four groups: 5-18 labeled as 0, 19-35 as 1, 36-59 as 2, and 60-88 as 3. Then the one-hot encodings of age and gender labels are processed through an MLP for encoding and fusion. The resulting embedding is then concatenated with the semantic embedding and passed to the decoder for generation.
To ensure a fair comparison, all experimental settings and hyperparameters (including seed, optimizer, epochs, scheduler, batch size, codebook size, etc.) strictly follow those in MoMask [27]. The weight coefficients $\alpha,\lambda$ in Equation 16 are set to $0.01,0.5$ empirically. Training is performed on a single RTX 4090 GPU, requiring 8 hours for the VQVAE and 26 hours for the generative transformer.
For the evaluation metrics, we follow the standard protocol [25] and report the following metrics: (1) Frechet Inception Distance (FID) [33], measuring the distance between feature distributions of generated and real motions; (2) R-Precision, the accuracy for motion-to-text retrieval in a batch of 32 samples; (3) Multimodal Distance (MM-Dist), measuring the Euclidean distance between the feature vectors of text and motion; (4) Diversity, which calculates the average Euclidean distance between features of 300 randomly sampled motion pairs; (5) Multimodality, assessing the diversity of generated motions for the given text. The feature extractor used for evaluation is pre-trained on the HumanAttr dataset, following the same method as in [25].
Methods R-Precision $\uparrow$ FID $\downarrow$ MM-Dist $\downarrow$
@@ 5 Experiments
Methods R-Precision $\uparrow$ FID $\downarrow$ MM-Dist $\downarrow$ Diversity $\rightarrow$ MModality $\uparrow$

### v2

@@ Abstract
Text-driven human motion generation has recently attracted considerable attention, allowing models to generate human motions based on textual descriptions. However, current methods neglect the influence of human attributes—such as age, gender, weight, and height—which are key factors shaping human motion patterns. This work represents a pilot exploration for bridging this gap. We conceptualize each motion as comprising both attribute information and action semantics, where textual descriptions align exclusively with action semantics. To achieve this, a new framework inspired by Structural Causal Models is proposed to decouple action semantics from human attributes, enabling text-to-semantics prediction and attribute-controlled generation. The resulting model is capable of generating attribute-aware motion aligned with the user’s text and attribute inputs. For evaluation, we introduce a comprehensive dataset containing attribute annotations for text-motion pairs, setting the first benchmark for attribute-aware motion generation. Extensive experiments validate our model’s effectiveness.
Introduction
@@
In recent years, text-driven human motion generation has gained significant attention (Guo et al. 2022a; Zhang et al. 2022; Chen et al. 2023) in both research community and industry owing to its potential applications in video games, filmmaking, virtual reality, and robotics. The goal is to generate realistic human motion sequences based on given textual descriptions. However, human motion patterns are significantly influenced by human attributes such as age, gender, weight, and height. As illustrated in Figure 1, individuals with different attributes demonstrate distinct motion patterns. For instance, the movements of an elderly person and a teenager differ substantially. Therefore, incorporating human attributes into motion generation is crucial for producing realistic and context-aware movements. Also, users may need to generate motions that correspond to specific human attributes, demanding text-to-motion generation controlled by attributes. Despite this, none of the existing methods have addressed this issue.
@@
Incorporating human attributes into text-driven motion generation faces two major challenges. The first is that a motion sequence consists of both action semantics (e.g., walking, running) and human attributes (e.g., gender, age), whereas textual descriptions usually focus solely on the action semantics aspect. Existing text-to-motion frameworks typically align text and motion in a shared space without distinguishing between semantic content and human attribute information, which may hinder the alignment. This necessitates a new framework that can separate action semantics from human attributes, ensuring that texts are accurately aligned with semantic content.
Figure 1: Data samples from HumanAttr dataset with text prompts and human attributes. Notice that the motion patterns of subjects with different attributes vary significantly.
@@
The second challenge lies in the absence of a large-scale text-to-motion dataset that includes annotations for a broad spectrum of human attributes. Existing datasets either lack human attribute annotations entirely (Guo et al. 2022a; Lin et al. 2024; Punnakkal et al. 2021; Liang et al. 2024b; Chen et al. 2024), encompass subjects with only a small range of distinct attributes (Plappert, Mandery, and Asfour 2016; Jang et al. 2020; Troje 2002; Ma et al. 2024; Kim et al. 2021), or have very limited scale (University and of Singapore ; Leightley et al. 2015). For example, the KIT dataset (Plappert, Mandery, and Asfour 2016) includes attribute annotations for 4k motion sequences, but 90% of the 55 subjects are aged between 18 and 45. This data gap impedes the development of attribute-aware motion generation models capable of generalizing to diverse textual descriptions and human attributes.
@@
This work represents a pilot exploration that incorporates human attributes to generate high-quality 3D human motion aligned with prompted text and attribute controls. To address the above-mentioned semantics-attribute coupling in human motions, we introduce AttrMoGen (Attribute-controlled Motion Generator). It comprises a Semantic-Attribute Decoupling VQVAE (Decoup-VQVAE) that extracts attribute-free semantic tokens from human motions, and a Semantics Generative Transformer that predicts these semantic tokens from textual prompts. To eliminate attribute information from the semantic tokens, we leverage Structural Causal Model (SCM) (Pearl and Mackenzie 2018), formulating the problem as causal factor disentanglement. In particular, action semantics $S$ are causal factors determining target tokens $Y$, while human attributes $A$ are non-causal for $Y$ but crucial for constructing raw motion $X$. The causal model is trained using a causal information bottleneck, which disentangles $S$ from $A$ by limiting their mutual information $I(S;A)$. Additionally, a bottleneck term $\lambda I(X;S)$ is used to restrict information flow from $X$ to $S$, ensuring that $S$ captures only the essential semantics. During inference, the Semantics Generative Transformer predicts semantic tokens from text input, which are combined with attributes input to generate the motion via the Decoup-VQVAE decoder.
@@
We compiled a comprehensive dataset called HumanAttr, which includes texts and attribute annotations for each motion sequence. HumanAttr integrates data from multiple sources, resulting in a collection of 18.2k motion sequences from 640 subjects with a wide range of attributes. Example data entries are presented in Figure 1. Each attribute label consists of the subject’s age and gender. A portion of the data ($\sim$ 74%) also contains weight and height information, but lacks age diversity thus are not utilized in the main experiments. An additional experiment on this subset that contains weight and height is included in supplementary materials for reference. Extensive experiments are conducted on the dataset, evaluating a range of current representative text-to-motion models, including our proposed AttrMoGen model. The results demonstrate its ability to generate realistic motion that aligns with texts and attributes.
@@
BMLmovi (Troje 2002) 86 1,801 161.8 [17, 33]
EKUT (Plappert, Mandery, and Asfour 2016) 4 348 30.1 [24, 44]
ETRI-Activity3D (Jang et al. 2020) 100 3,727 691.8 [21, 88]
ETRI-LivingLab (Kim et al. 2021) 50 600 160.5 [64, 80]
K3Da (Leightley et al. 2015) 54 236 34.5 [18, 81]
Kinder-Gator (Aloba et al. 2018) 20 362 34.9 [5, 32]
KIT (Plappert, Mandery, and Asfour 2016) 55 4,231 463.1 [15, 55]
Nymeria (Ma et al. 2024) 264 6,850 552.4 [18, 50]
SFU (University and of Singapore ) 7 44 6.4 [18, 30]
@@
Table 1: Statistics of the collected sub-datasets. The statistics are gathered after data filtering.
Related Work
Human Motion Generation. Generating realistic human motion that aligns with specific control signals has been a longstanding problem. A prominent focus is text-to-motion generation. Early works (Tevet et al. 2022a; Petrovich, Black, and Varol 2022, 2023; Yu, Tanaka, and Fujiwara 2024) primarily focused on establishing a joint embedding space for motion sequences and texts. Recent development has been dominated by autoregressive models and diffusion models. The former (Guo et al. 2022a; Huang et al. 2025; Lu et al. 2023) represent motions with discrete tokens and implement text-to-motion generation via next token prediction, while diffusion models (Jin et al. 2024; Yuan et al. 2023; Zhang et al. 2024b; Wang et al. 2023; Karunratanakul et al. 2023; Chen et al. 2023; Xie et al. 2024b, a; Liang et al. 2024a; Huang et al. 2024) perform denoising in the motion space or latent space to generate natural and smooth human motions. Recently, some studies (Zhong et al. 2024; Guo et al. 2024b; Kim et al. 2024; Song et al. 2024) explore style-based motion models, emphasizing how an action is performed (e.g., proud, depressed, angry). Yet, style refers to subjective intent behind an action, while attributes determine the objective inherent biomechanics (e.g., child and the elderly naturally differ in stride length, joint range, and expressiveness).
Text-motion Datasets. With the proliferation of mocap-based databases (CMU Graphics Lab 2000; Shahroudy et al. 2016; Liu et al. 2019; Mandery et al. 2015; Mahmood et al. 2019), numerous text-motion datasets have been proposed. Conventional ones like KIT and HumanML3D provide motion with corresponding sequence-level textual descriptions, primarily developed for text-to-motion task. Subsequent works have advanced the field by incorporating multi-person scenarios (Liang et al. 2024b; Fang et al. 2024), addressing long motion sequences (Li et al. 2023; Han et al. 2024), enhancing the quality of text descriptions (Tang et al. 2023; Chen et al. 2024), adding full-body text descriptions (Lin et al. 2024), and providing frame-level or segment-level annotations (Punnakkal et al. 2021; Wang, Kang, and Mu 2024). Several motion datasets (University and of Singapore ; Plappert, Mandery, and Asfour 2016; Troje 2002; Ma et al. 2024; Jang et al. 2020; Kim et al. 2021) provide the attribute information of each subject, yet are limited in data scale or attribute range. For instance, 90% of the subjects in KIT are aged between 18 and 45. This indicates the lack of large-scale collections that encompass a wide range of human attributes.
@@
The HumanAttr Dataset
Dataset Collection. Existing benchmark datasets for human motion generation such as HumanML3D and KIT primarily consist of motion sequences and corresponding text descriptions, lacking the annotations for human attributes. To address this gap, we compile a comprehensive dataset that includes data entries of (motion, text, attribute). The attribute labels contain age and gender for each data entry, and a portion of the data ($\sim$74%) also have weight and height information. The dataset, named HumanAttr, integrates data from various sources, annotated with diverse human attributes and textual descriptions. As illustrated in Table 1, HumanAttr consists of nine sub-datasets that encompass a variety of subjects with varying attributes. In total, it contains 18.2k motion sequences from 640 subjects, with ages ranging from 5 to 88 years and balanced gender. Figure 7 presents the distribution of age, gender, motion duration, and text length within the dataset.
Dataset Processing. This section provides a brief introduction to the pipeline for processing HumanAttr. Initially, data from different sources are converted into a 3D coordinates representation of $L\times V\times 3$, where $L$ denotes the sequence length and $V$ is the number of markers. Since the number of markers varies across different data sources, we adopt the method from (Li et al. 2024) and use gradient descent optimization to fit the marker data into unified parameterized SMPL model (Loper et al. 2023). To make the dataset compatible with previous work in the field, processing protocol from HumanML3D (Guo et al. 2022a) is then applied, resulting in a 263-dimensional input vector.
However, some of the motion data exhibited serious jittering issues caused by the Kinect system. To address this, additional processing is applied to specific sub-datasets. Motions with excessive jitter or outlier coordinates are discarded, and a 1-D Gaussian denoising is applied to enhance motion quality. The dataset is further augmented by mirroring the motion data following the approach in HumanML3D to increase the training samples and improve model generalization. For text annotations, all sub-datasets include either textual descriptions or action category labels. Action labels are expanded into full text descriptions using the information from the original database, ensuring each motion entry has a meaningful text.
Our Proposed Method
To tackle this challenge, we propose an architecture composed of two main components. The overall architecture is illustrated in Figure 4. The first component is a Semantic-attribute Decoupling VQVAE (Decoup-VQVAE), inspired by Structural Causal Model (SCM) (Pearl and Mackenzie 2018). Specifically, the encoder removes attribute information from the raw motion with the help of a causal information bottleneck, resulting in attribute-free semantic tokens via vector quantization. The decoder can then reconstruct the original motion using both the semantic tokens and attribute labels. The second component is a Semantics Generative Transformer, which predicts semantic tokens from text by a generative transformer. During inference, semantic tokens are generated from textual input and combined with user-defined attributes to generate attribute-aware human motion.
Figure 3: Structural Causal Model for our Decoup-VQVAE. Our objective is to learn an encoder capable of decoupling the $Y$-causative action semantics $S$ from raw motion $X$.
@@
Semantic-attribute Decoupling VQVAE
To obtain attribute-free semantic tokens, we employed Structural Causal Model (SCM) to formulate the problem as causal factor disentanglement. Specifically, let $X$ represent the raw motion and $Y$ represent the desired semantic tokens. The action semantics $S$ are the causal factors that determine $Y$. Human attributes, denoted as $A$, are non-causal factors that are independent of $Y$, but essential in constructing the raw motion $X$. This leads to our causal formulation, with the SCM illustrated in Figure 3. Our aim is to learn an encoder $S=f(X)$ that can decouple the $Y$-causative action semantics $S$ from raw motion $X$, and a decoder $\hat{X}=g(S,A)$ that can reconstruct the motion $X$ from $S,A$. Ideally, $f$ should clearly separate $S$ from $A$, retaining only the essential semantic information of the raw motion. Inspired by (Zhang et al. 2024a; Hua et al. 2022), we introduce a causal information bottleneck (CIB) to tackle this challenge. Formally, The CIB objective function is defined as
@@
where $I(;)$ denotes the mutual information and $\lambda$ is the bottleneck weight. (Zhang et al. 2024a) proves that maximizing the CIB objective is equivalent to maximizing the mutual information $I(X,Y;S,A)$ between the latent factors $S$ and $A$ and the variables $X$ and $Y$ with an information bottleneck. The role of each term and our customized implementation will be demonstrated below.
@@
The term I(S;A)-I(S;A) restricts the mutual information between $S$ and $A$, which plays a crucial role in disentangling the causal factor $S$ from the non-causal factor $A$. In our case, note that the mutual information has the following upper bound:
@@
The information bottleneck term $\lambda I(X;S)$ limits the information flow from $X$ to $S$, ensuring that $S$ contains only essential semantics without excessive irrelevant information, fostering a robust and compact representation (Tishby, Pereira, and Bialek 2000). The mutual information $I(X;S)$ can be measured using the Kullback-Leibler (KL) divergence:
@@
where $A^{-}$ is the counterfactual attribute obtained by randomizing the original attribute $A$. Suppose $\boldsymbol{X}$ is a batch of samples and $\boldsymbol{X}^{-}$ is their corresponding counterfactuals. As the encoder $S=f(X)$ serves as an estimator for the distribution $p(S|X)$, the KL-divergence in Equation 10 can be minimized by aligning $f(\boldsymbol{X})$ and $f(\boldsymbol{X}^{-})$. To achieve this, we define the following similarity matrix analogous to (Lv et al. 2022):
@@
The remaining terms, $I(X;S,A)+I(Y;S)$, ensure that the information contained in $S$ and $A$ are sufficient to reconstruct $X$, while also guarantee that $S$ is informative enough to deduce $Y$. Note the following lower bound derived by (Zhang et al. 2024a):
@@
The second term $I(Y;S)$ ensures the deduction from the causal factor $S$ to the target semantic token $Y$, which is, by definition, the quantization process from the semantic embedding to the codebook index. Note that I(Y;S)=H(Y)H(Y|S)I(Y;S)=H(Y)-H(Y|S), where $H(Y)$ reflects the full exploitation of the codebook and H(Y|S)-H(Y|S) demands precise indexing from $S$ to $Y$. This can be addressed by the loss $\mathcal{L}_{commit}+\mathcal{L}_{embed}$ from VQVAE (Van Den Oord, Vinyals et al. 2017). In summary, the overall loss function is formulated as
@@
Semantics Generative Transformer
To predict semantic tokens from text, a Semantics Generative Transformer is employed, whose architecture can be either a GPT-like autoregressive model (Zhang et al. 2023; Lu et al. 2023) or a masked transformer (Guo et al. 2024a). We adopt MoMask (Guo et al. 2024a) due to its strong performance and efficiency. During training, the model aims to predict randomly masked semantic tokens conditioned on the CLIP (Radford et al. 2021) features of the textual input, following a BERT-like scheme (Devlin 2018). At inference time, semantic tokens are first predicted from the text using the Semantics Generative Transformer. These tokens, combined with attribute inputs (e.g., age, gender), are then used to generate the motion via the Decoup-VQVAE decoder. Figure 4 presents a schematic illustration.
Methods R-Precision $\uparrow$ FID $\downarrow$ $M$-Dist $\downarrow$
@@
Methods R-Precision $\uparrow$ FID $\downarrow$ $M$-Dist $\downarrow$ Diversity $\rightarrow$ MModality $\uparrow$